# Language Model Planning from an Information Theoretic Perspective

### v1

@@ 2 Method
To estimate MI between the high-level representations of hidden state blocks, $G_{\mathcal{S}}$, we need to find a mapping from $G_{\mathcal{S}}$ to the discrete codes. The size of target layer-token indices $\mathcal{S}$ and its corresponding activation outputs $G$ may vary depending on the analysis. For instance, when examining all block outputs across the prefix tokens, $\mathcal{S}$ spans every layer and grows with the prefix length $T$. Since neural networks are good at learning representations, we employ an encoder-decoder neural network architecture: the encoder maps the potentially long and variable-sized $G$ into a compact representation, and the decoder aims at recovering $G$ from it. Leveraging the effectiveness in handling variable-length inputs and learning rich representations [31], we adopt transformer-based encoder and decoder models. Specifically, we employ Vector-Quantized Variational Autoencoders (VQ-VAEs; Van Den Oord et al. 57) framework. VQ-VAE is a framework consisting of an encoder-decoder network with a trainable embedding codebook between them to discretize the encoder’s output. Its training objective consists of a reconstruction loss between the decoder’s output and the input signal, supplemented with auxiliary terms that encourage effective utilization and learning of the codebook embeddings. Overall, a VQ-VAE acts as a quantizer that extracts compact representations of the input by mapping to a codebook index. In addition to the vanilla VQ-VAE’s training objective, we introduce a cosine similarity loss across codebook embeddings to encourage representations assigned to different codes to diverge so that they are not only compact but also discriminative. [^2] In short, our VQ-VAE setting with the additional cosine loss combines the regularizing and contrastive approaches for representation learning, e.g., [11, 17] This enables embeddings to spread out in the latent space and ensures that codes capture distinct features of the data making them informative and discriminative. We refer reader to App. A.1 for the details of the implementation.
@@
The trained encoder and the codebook of VQ-VAE are used to obtain coarse representations of transformer computations. Given a set $G$ of block outputs corresponding to indices in $\mathcal{S}$, we map $G$ to a discrete code in $[K]$, where $K$ is the size of the codebook. Instead of fine-grained activations, we work with these high-level representation codes that serve as compact summary of complex internal computations in $G$. Importantly, we treat these codes not as semantically meaningful entities, but as representational summaries that allow us to find information relations. Therefore, the estimated nMI depends not only on the underlying relationship between the hidden state themselves, but also on how they are encoded to high-level codes. Our use of VQ-VAE ensures that the learned representations are nontrivial since the decoder achieves meaningful reconstructions, and the cosine similarity penalty enforces diversity among codebook embeddings, which is validated by the reconstruction quality and codebook similarity results presented in App. A.1. Furthermore, we validate that our framework can find a closer relation to a known underlying distribution, in a representative experimental setting. Due to space limits, we refer the reader to App. A.3 for the details and results of the validation experiment.
@@ 3 Analysis of planning capability
We leverage our method to investigate three core aspects of the planning capabilities of LMs and their underlying computations: (3.1) how much a model plans about future tokens before producing its immediate prediction, (3.2) to what extent the model internally considers alternative correct continuations, and (3.3) to what extent the model relies on earlier computations within the context window when deciding the next tokens. In our analyses, we use architecture-matched GPT-3 Small [9] models employing rotary position embeddings [52]. We report results from multiple seeds.
@@
In our analysis, we examine two variants of LM training objectives: $\mathcal{M}_{\text{NTP}}$ is trained with the standard next-token prediction (NTP) loss ($\mathcal{L}_{\text{NTP}}$), which optimizes conditional likelihood of the immediate next token, and $\mathcal{M}_{\text{MTP}}$ is trained with the multi-token prediction (MTP) loss ($\mathcal{L}_{\text{MTP}}$), which encourages models to align their hidden states with predictions across multiple forthcoming tokens. Following [18] for MTP implementation, we use shared transformer layers and separate heads for the next tokens during training. During inference, both models use standard autoregressive next-token generation. Intuitively, NTP prioritizes local token-by-token consistency, whereas MTP incentivizes models to form short-horizon plans and reduces strictly myopic behavior [39, 5]. Both losses are described in detail in App. B.1.
@@
Our experiments span three datasets chosen to cover distinct aspects of planning. (i) A context-free grammar (CFG) dataset [4] emphasizes token-level rules and local coherence. Tokens in CFG carry intrinsic meaning, much like words in natural language, and are governed by both local and global syntactic constraints. The model’s task is to learn the grammar and predict valid next tokens that obey its rules. As a controlled substitute for natural language, CFG allows us to adjust vocabulary size and task difficulty, providing a clean environment for scientific study. (ii) A path-finding (PF) task (see Figure 1, right for illustration) requires models to generate valid paths between a fixed start and goal node in graphs with varying sizes and edge structures. The node vocabulary consists of 28 unique tokens, each representing a node identity. Unlike in CFG, these tokens do not carry intrinsic semantics and the meaning of a node token arises only through the set of edges it participates in. To construct language-model inputs, we linearize the graph by writing edges as pairs of adjacent node tokens separated by commas in a random order to prevent the model from exploiting positional shortcuts (e.g., “$u$ $v$, $r$ $t$, …”). This edge-list representation ensures that the model must infer connectivity and graph structure rather than rely on token identity alone. Solving PF thus requires composing multiple reasoning steps over these relational inputs, analogous to chaining lemmas to form a proof, introducing a natural flavor of long-horizon planning. We construct two variants, PF-Short and PF-Long, which require finding paths of length 4 and 6 (including start and goal), respectively. (iii) A natural language dataset, OpenWebText [19], reflects real-world distributional modeling. Here the model predicts continuations consistent with natural text. These datasets cover a spectrum from low-level symbolic structure to high-level naturalistic text. We refer the reader to App. B.2 for details.
Figure 1: The visualizations of experimental settings. Left: History in the plan experiment (§ 3.3). Middle: Horizon of the plan and branches in the plan experiment (§ 3.1 & § 3.2). The red, blue, and green colors indicate target variables used in our analysis in left and middle panels. Right: An illustration of a simplified sample in PF task. Graphs and token numbers are randomly generated except for the start (1) and goal node (2). A sample prompt is “19 23 , 13 2 , 11 8 , 4 17 , 1 23 , 9 7, 2 8, 17 1 , 13 2, 14 23, 23 7, 1 23 :" and correct responses are “1 11 8 2" or “1 17 13 2".
@@ 3.1 Horizon of the plan
where $H$ encodes the entirety of the model’s prefix computation that has an effect on future token generations. Training a separate VQ-VAE over $H$, we get the LM’s pre-output computation summary, and denote its high-level code as $Z_{\scriptstyle 1:T}^{1:L-1}$. $H$ and $h_{\scriptscriptstyle T+\tau}^{\scriptscriptstyle L}$} are respectively shown in red and blue colors in Figure 1 (left). We address our question by comparing how much the prefix computation’s summary $Z_{\scriptstyle 1:T}^{1:L-1}$ tells us about the model’s decision state at the generated tokens. We measure the nMI between $Z_{\scriptstyle 1:T}^{1:L-1}$ and $Z_{\scriptstyle T+\tau}^{\scriptstyle L}$ for $\tau\geq$0 as defined in Eq. 2 with normalization $\mathcal{I}_{\max}=\max\{\mathcal{I}(Z_{\scriptstyle 1:T}^{1:L-1};Z^{\scriptstyle L}_{\scriptstyle T+n})\}_{n}$. If this ratio remains significantly above zero for large $\tau$, then the pre-output computation $Z_{\scriptstyle 1:T}^{1:L-1}$ carries information on the strategy to produce $\widehat{x}_{T+\tau}$, indicating the model’s initial prefix processing is not merely myopic but encodes forward-looking structure that persists into later generations. Conversely, if the ratio quickly decays to zero, then prefix computations primarily support only the immediate next prediction, with little evidence of planning ahead.
@@
We generate sentences of length $[16,67]$ from CFG rules, and for each sample select a prefix length $T\in[8,24]$ uniformly at random before, with the model generating the remaining continuation. For $\mathcal{M}_{\text{NTP}}$ and $\mathcal{M}_{\text{MTP}}$, Figure 2(a) reports the nMI between $Z_{\scriptstyle 1:T}^{1:L-1}$ and $Z_{\scriptstyle T+\tau}^{\scriptstyle L}$ across $\tau\geq 0$. Because CFG tokens have intrinsic meaning and are significantly governed by local syntactic constraints, it is expected that prefix computations correlate most strongly with the first few generated tokens. Indeed, $Z_{\scriptstyle 1:T}^{1:L-1}$ retains the highest information about the decision state for $\tau=0$, and by $\tau=10$ the nMI drops to roughly one-fifth of its initial value. This indicates that the pre-output computation encodes a short-horizon plan tied mainly to the next few tokens. $\mathcal{M}_{\text{MTP}}$ exhibits a similar pattern, with a slightly slower decay for $\tau<10$, suggesting that the MTP loss does not substantially extend the horizon of pre-output computation in a model well-trained on CFG.
@@
Figure 2: nMI results between the prefix summary codes and the last hidden state codes of generated tokens for CFG (a) and PF (b) tasks. nMI decays fast in CFG, showing short and local planning, while PF maintains or even increases nMI beyond $\tau=1$, reflecting planning for later steps.
@@
On PF-Short and PF-Long, we trained models on NTP and MTP losses. Because the start and end nodes are fixed tokens, the challenge lies in predicting the correct intermediate nodes ($\tau=1,2$ for PF-Short and $\tau=1,2,3,4$ for PF-Long) that connect them. We therefore focus our analysis on the intermediate positions and compute nMI as defined in Eq. 2. Figure 2(b) shows nMI across the generated path as well as the accuracy of correctly finding the whole path. For PF-Short, we see similar nMI trends across generated tokens for both $\mathcal{M}_{\text{MTP}}$ and $\mathcal{M}_{\text{NTP}}$, which both attain high accuracy. Interestingly, prefix computations encode more information about the second intermediate node than about the first. A likely explanation is that there is less uncertainty about the final token given the edge information and that the model spares more pre-output computation to find the last token. This resembles a strategy to work backwards from the goal, as a human solving this task might do. Similarly, for PF-Long, in Figure 2(b), both $\mathcal{M}_{\text{NTP}}$ and $\mathcal{M}_{\text{MTP}}$ prefix computations encode a comparatively high amount of MI about the future tokens, unlike to the steady decay seen in CFG. This suggests that pre-output computations are not limited to the next token but also embed plans for upcoming positions, reflecting deliberate allocation of capacity toward future outputs. This finding is aligned with previous research on training LMs to exploit this reverse-solving approach, which has been shown to improve performance [5]. Taken together, these results indicate that LMs trained with both NTP and MTP loss can exhibit non-myopic behavior when trained on tasks that demand it. Lastly, in the PF-Long experiment, we observe that the nMI of $\mathcal{M}_{\text{MTP}}$ is slightly more uniform across $\tau$ than that of $\mathcal{M}_{\text{NTP}}$, which is a likely explanation for its higher accuracy by better predicting the earlier (harder) tokens along the generated path.
@@ 3.3 History in the plan
To what extent do LMs rely on earlier computations when generating new tokens?
To answer the question, we use an 8-layer decoder-only Transformer trained on OpenWebText with NTP objective. We estimate the nMI between the codes of the model’s decision state for the generated tokens and of blocks of prefix hidden state computations within the LM. Similar to § 3.1, we obtain a summary of the strategy employed for token $\widehat{x}_{T+\tau+1}$, denoted $Z_{\scriptstyle T+\tau}^{\scriptstyle L}$ with a VQ-VAE trained over the last layer of hidden states. We also partition each layer’s prefix hidden states into contiguous, non-overlapping blocks of length 16 and train a VQ-VAE to acquire the code $\mathbf{B}_{k}^{\ell}$ where the block index $k\in\{1,\dots,12\}$ denotes the $k^{\text{th}}$ block from the end of the prefix (i.e., $k=1$ corresponds to the most recent 16 time steps before generation). In Figure 1 (middle), a sample block and a sample last hidden state are illustrated. We quantify the dependence between the codes of each computation block and the decision state by measuring nMI between $\mathbf{B}_{k}^{\ell}$ and $Z_{\scriptstyle T+\tau}^{\scriptstyle L}$} across all layers $\ell$ and blocks $k$, as defined in Eq. 2, with $\mathcal{I}_{\max}=\max\{\mathcal{I}(\mathbf{B}_{k}^{\ell};Z^{L}_{T+n})\}_{\ell,k,n}$. Heatmaps in Figure 3 (left and middle) show the results for $\tau$= 0 and 1 (see App. B.5.1 for $\tau>1$).
We observe that last layer computations retain the most information about the decision state of both immediate and future tokens which is aligned with the results of [41] and consistent with the design choice of assigning longer attention spans to higher layers [53, 6]. Furthermore, along the block axis, we observe a clear recency effect across all layers: the codes of the most recent blocks, i.e., final time steps of the prefix, computations ($Z_{1}$), exhibit the highest nMI (Eq. 2) with the decision state of the generated tokens, and the nMI decays over earlier blocks in the prefix. Although analysis along both axes indicates that the LM primarily relies on the most recent computations, we still observe appreciable nMI in lower layers (small $\ell$) and earlier blocks (large $k$). This suggests that LMs retain information from earlier parts of the prefix when generating new tokens, instead of relying only on the most recent computations.
[…]

### v2

@@ 2 Method
To estimate MI between the high-level representations of hidden state blocks, $G_{\mathcal{S}}$, we need to find a mapping from $G_{\mathcal{S}}$ to the discrete codes, as depicted in Step 1 of Fig. 1. The size of target index set $\mathcal{S}$ and its corresponding activation outputs $G$ may vary depending on the analysis. For instance, when examining all block outputs across the prefix tokens, $\mathcal{S}$ spans every layer and grows with the prefix length $T$. Since neural networks are effective at learning representations, we employ an encoder-decoder neural network architecture: the encoder maps the potentially long and variable-sized $G_{\mathcal{S}}$ to a compact representation, and the decoder aims to recover $G_{\mathcal{S}}$ from it. Leveraging the effectiveness in handling variable-length inputs and learning rich representations (Lin et al., 2022), we adopt transformer-based encoder and decoder models. Specifically, we employ Vector-Quantized Variational Autoencoder (VQ-VAEs; Van Den Oord et al. 2017). VQ-VAE is a framework consisting of an encoder-decoder network with a trainable embedding codebook between them to discretize the encoder’s output.
Concretely, for each family of index sets $\mathcal{S}$ used in our analyses, we train a separate single VQ-VAE consisting of an encoder $E$, a codebook $\{e_{k}\}_{k=1}^{K}\subset\mathbb{R}^{d_{e}}$, and a decoder $D$. Given a block $G_{\mathcal{S}}$, the encoder produces a latent vector $r_{\mathcal{S}}=E(G_{\mathcal{S}})\in\mathbb{R}^{d_{e}}$. We then quantize by nearest neighbor in the codebook,
$k^{\star}\;=\;\arg\min_{k\in[K]}\|r_{\mathcal{S}}-e_{k}\|_{2}^{2},\qquad Z_{\mathcal{S}}\equiv k^{\star},\qquad\tilde{r}_{\mathcal{S}}\equiv e_{k^{\star}},$
and reconstruct $\widehat{G}_{\mathcal{S}}=D(\tilde{r}_{\mathcal{S}})$. During the training of VQ-VAE model, gradients are propagated through the quantization step using the straight-through estimator (Van Den Oord et al., 2017). The overall training objective combines reconstruction, standard vector-quantization, and two regularization terms that promote informative and well-used codes:
$\mathcal{L}=\mathcal{L}_{\text{rec}}+\lambda_{q}\,\mathcal{L}_{\text{vq}}+\lambda_{\text{cos}}\,\mathcal{L}_{\text{cos}}+\lambda_{\text{ent}}\,\mathcal{L}_{\text{ent}}.\vskip-3.00003pt$ (3)
In Eq. 3, $\mathcal{L}_{\text{rec}}$ is a reconstruction loss that encourages $\widehat{G}_{\mathcal{S}}$ to match $G_{\mathcal{S}}$, $\mathcal{L}_{\text{vq}}$ is the quantization and commitment loss as in Van Den Oord et al. (2017) tying encoder outputs to discrete codes. In addition to vanilla VQ-VAE’s training objective we introduce $\mathcal{L}_{\text{cos}}$, a cosine-similarity penalty that pushes codebook embeddings $\{e_{k}\}$ to be diverse, and $\mathcal{L}_{\text{ent}}$, an entropy term that discourages collapse to a small subset of codes. These losses encourage representations assigned to different codes to diverge so that they are not only compact but also discriminative. [^2] In short, our VQ-VAE setting with the additional cosine loss combines the regularizing and contrastive approaches for representation learning, e.g., (Chen et al., 2020; Gao et al., 2021) This enables embeddings to spread out in the latent space and ensures that codes capture distinct features of the data making them informative and discriminative. Details about the formulas and implementation are given in App. A.1 and App. A.2.
Overview of the pipeline.
@@
Fig. 1 summarizes the procedure. For a frozen pretrained LM $\mathcal{M}$, we run prefixes ${\mathbf{x}}_{1:T}$, select block index sets $\mathcal{S}$, and extract the corresponding hidden-state blocks $G_{\mathcal{S}}$. For each family of blocks, we then train a VQ-VAE so that its encoder maps each $G_{\mathcal{S}}$ to a discrete code $Z_{\mathcal{S}}\in[K]$, where $K$ is the codebook size (Step 1). Instead of fine-grained activations, we work with these high-level representation codes that serve as compact summary of complex internal computations in $G$. Importantly, we treat these codes not as semantically meaningful entities, but as representational summaries that allow us to find information relations. Therefore, the estimated mutual information depends not only on the underlying relationship between the hidden state themselves, but also on how they are encoded to high-level codes. Our use of VQ-VAE ensures that the learned representations are nontrivial since the decoder achieves meaningful reconstructions, and the cosine similarity penalty enforces diversity among codebook embeddings, which is validated by the reconstruction quality and codebook similarity results presented in App. A.1. In Step 2, we apply the trained encoder and codebook to the full dataset to obtain codes $Z_{A}$ and $Z_{B}$ for two compared block collections $G_{A}$ and $G_{B}$, count their co-occurrences. Then, we using the empirical joint and marginal distributions, in Step 3, we find $I(Z_{A};Z_{B})$ and its normalized variant $\text{nMI}(Z_{A};Z_{B})$, used as measures of how information is shared between different parts of the model’s computation. Furthermore, we validate that our framework can find a closer relation to a known underlying distribution, in a representative experimental setting. Due to space limits, we refer the reader to App. A.4 for the details and results of the validation experiment. Experiments discussing probing-based method (Xu et al., ) and a diagnostic illustrating why probing can have confounding effects are presented in App. D.2D.5.
@@ 3 Analysis of planning capability
We leverage our method to investigate two core planning-related aspects—(3.1) how much a model’s prefix computation is forward-looking about future tokens, and (3.2) whether it preserves alternative correct continuations—and, as an application of the same framework, (3.3) how decision-relevant information is encoded across layers and earlier prefix blocks when producing next-token decisions. In our analyses, we use architecture-matched GPT-3 Small (Brown, 2020) models employing rotary position embeddings (Su et al., 2024). We report results from multiple seeds.
@@
In our analysis, we examine two variants of LM training objectives: $\mathcal{M}_{\text{NTP}}$ is trained with the standard next-token prediction (NTP) loss ($\mathcal{L}_{\text{NTP}}$), which optimizes conditional likelihood of the immediate next token, and $\mathcal{M}_{\text{MTP}}$ is trained with the multi-token prediction (MTP) loss ($\mathcal{L}_{\text{MTP}}$), which encourages models to align their hidden states with predictions across multiple forthcoming tokens. Following Gloeckle et al. (2024) for MTP implementation, we use shared transformer layers and separate heads for the next tokens during training. During inference, both models use standard autoregressive next-token generation. Intuitively, NTP prioritizes local token-by-token consistency, whereas MTP incentivizes models to form short-horizon plans and reduces strictly myopic behavior (Nagarajan et al., 2025; Bachmann and Nagarajan, 2024). Both losses are described in detail in App. B.1.
@@
Our experiments span three datasets chosen to cover distinct aspects of planning. (i) A context-free grammar (CFG) dataset (Allen-Zhu, 2024) emphasizes token-level rules and local coherence. Tokens in CFG carry intrinsic meaning, much like words in natural language, and are governed by both local and global syntactic constraints. The model’s task is to learn the grammar and predict valid next tokens that obey its rules. As a controlled substitute for natural language, CFG allows us to adjust vocabulary size and task difficulty, providing a clean environment for scientific study. (ii) A path-finding (PF) task (see Figure 2, right for illustration) requires models to generate valid paths between a fixed start and goal node in graphs with varying sizes and edge structures. The node vocabulary consists of 28 unique tokens, each representing a node identity. Unlike in CFG, these tokens do not carry intrinsic semantics and the meaning of a node token arises only through the set of edges it participates in. To construct language-model inputs, we linearize the graph by writing edges as pairs of adjacent node tokens separated by commas in a random order to prevent the model from exploiting positional shortcuts (e.g., “$u$ $v$, $r$ $t$, …”). This edge-list representation ensures that the model must infer connectivity and graph structure rather than rely on token identity alone. Solving PF thus requires composing multiple reasoning steps over these relational inputs, analogous to chaining lemmas to form a proof, introducing a natural flavor of long-horizon planning. We construct two variants, PF-Short and PF-Long, which require finding paths of length 4 and 6 (including start and goal), respectively. (iii) A natural language dataset, OpenWebText (Gokaslan et al., 2019), reflects real-world distributional modeling. Here the model predicts continuations consistent with natural text. These datasets cover a spectrum from low-level symbolic structure to high-level naturalistic text. We refer the reader to App. B.2 for details.
Figure 2: The visualizations of experimental settings. Left: Horizon of the plan and branches in the plan experiment (§ 3.1 & § 3.2). Middle: Information in the computational history experiment (§ 3.3). Colors in the left and middle panels denote the target variables. Right: An illustration of a simplified sample in PF task. Graphs and token numbers are randomly generated except for the start (1) and goal node (2). A sample prompt is “19 23 , 13 2 , 11 8 , 4 17 , 1 23 , 9 7, 2 8, 17 1 , 13 2 , 14 23 , 23 7 , 1 23 :” and correct responses are “1 11 8 2” or “1 17 13 2”.
@@ 3.1 Horizon of the plan
where $H$ encodes the entirety of the model’s prefix computation that has an effect on future token generations. Training a separate VQ-VAE over $H$, we get the LM’s pre-output computation summary, and denote its high-level code as $Z_{\scriptstyle 1:T}^{1:L-1}$. $H$ and $h_{\scriptscriptstyle T+\tau}^{\scriptscriptstyle L}$} are respectively shown in red and blue colors in Figure 2 (left). We address our question by comparing how much the prefix computation’s summary $Z_{\scriptstyle 1:T}^{1:L-1}$ tells us about the model’s decision state at the generated tokens. We measure the nMI between $Z_{\scriptstyle 1:T}^{1:L-1}$ and $Z_{\scriptstyle T+\tau}^{\scriptstyle L}$ for $\tau\geq$0 as defined in Eq. 2 with normalization $\mathcal{I}_{\max}=\max\{\mathcal{I}(Z_{\scriptstyle 1:T}^{1:L-1};Z^{\scriptstyle L}_{\scriptstyle T+n})\}_{n}$. If this ratio remains significantly above zero for large $\tau$, then the pre-output computation $Z_{\scriptstyle 1:T}^{1:L-1}$ carries information on the strategy to produce $\widehat{x}_{T+\tau}$, indicating the model’s initial prefix processing is not merely myopic but encodes forward-looking structure that persists into later generations. Conversely, if the ratio quickly decays to zero, then prefix computations primarily support only the immediate next prediction, with little evidence of long-horizon dependence between prefix computation and later decision states.
@@
We generate sentences of length $[16,67]$ from CFG rules, and for each sample select a prefix length $T\in[8,24]$ uniformly at random before, with the model generating the remaining continuation. For $\mathcal{M}_{\text{NTP}}$ and $\mathcal{M}_{\text{MTP}}$, Figure 3(a) reports the nMI between $Z_{\scriptstyle 1:T}^{1:L-1}$ and $Z_{\scriptstyle T+\tau}^{\scriptstyle L}$ across $\tau\geq 0$. Because CFG tokens have intrinsic meaning and are significantly governed by local syntactic constraints, it is expected that prefix computations correlate most strongly with the first few generated tokens. Indeed, $Z_{\scriptstyle 1:T}^{1:L-1}$ retains the highest information about the decision state for $\tau=0$, and by $\tau=10$ the nMI drops to roughly one-fifth of its initial value. This indicates that the pre-output computation encodes a short-horizon plan tied mainly to the next few tokens. We observe the same rapid decay of nMI when scaling the LM to 0.30.3B parameters (App. D.1. $\mathcal{M}_{\text{MTP}}$ exhibits a similar pattern, with a slightly slower decay for $\tau<10$, suggesting that the MTP loss does not substantially extend the horizon of pre-output computation in a model well-trained on CFG.
@@
Figure 3: nMI results between the prefix summary codes and the last hidden state codes of generated tokens for CFG (a) and PF (b) tasks. nMI decays fast in CFG, consistent with short-range dependence, while PF maintains or even increases nMI beyond $\tau=1$, consistent with longer-horizon predictive dependence of prefix computation on later decision states.
@@
On PF-Short and PF-Long, we trained models on NTP and MTP losses. Because the start and end nodes are fixed tokens, the challenge lies in predicting the correct intermediate nodes ($\tau=1,2$ for PF-Short and $\tau=1,2,3,4$ for PF-Long) that connect them. We therefore focus our analysis on the intermediate positions and compute nMI as defined in Eq. 2. Figure 3(b) shows nMI across the generated path as well as the accuracy of correctly finding the whole path. For PF-Short, we see similar nMI trends across generated tokens for both $\mathcal{M}_{\text{MTP}}$ and $\mathcal{M}_{\text{NTP}}$, which both attain high accuracy. Interestingly, prefix computations encode more information about the second intermediate node than about the first. A likely explanation is that there is less uncertainty about the final token given the edge information and that the model spares more pre-output computation to find the last token. This resembles a strategy to work backwards from the goal, as a human solving this task might do. Similarly, for PF-Long, in Figure 3(b), both $\mathcal{M}_{\text{NTP}}$ and $\mathcal{M}_{\text{MTP}}$ prefix computations encode a comparatively high amount of MI about the future tokens, unlike to the steady decay seen in CFG. This suggests that pre-output computations are not limited to the next token but also embed plans for upcoming positions, reflecting deliberate allocation of capacity toward future outputs. This finding is aligned with previous research on training LMs to exploit this reverse-solving approach, which has been shown to improve performance (Bachmann and Nagarajan, 2024). Taken together, these results indicate that LMs trained with both NTP and MTP loss can exhibit non-myopic behavior when trained on tasks that demand it. Lastly, in the PF-Long experiment, we observe that the nMI of $\mathcal{M}_{\text{MTP}}$ is slightly more uniform across $\tau$ than that of $\mathcal{M}_{\text{NTP}}$, which is a likely explanation for its higher accuracy by better predicting the earlier (harder) tokens along the generated path.
@@ 3.3 Information in the computational history
Which earlier layers and prefix blocks remain informative about the next-token decision state?
We view this primarily as a diagnostic. It does not tell us how to improve the model, but it quantifies how concentrated next-token decision information is in recent vs. earlier computation, which can help design choices that alter attention span or depth-wise context allocation. To answer the question, we use an 8-layer decoder-only Transformer trained on OpenWebText with NTP objective. We estimate the nMI between the codes of the model’s decision state for the generated tokens and of blocks of prefix hidden state computations within the LM. Similar to § 3.1, we obtain a summary of the strategy employed for token $\widehat{x}_{T+\tau+1}$, denoted $Z_{\scriptstyle T+\tau}^{\scriptstyle L}$ with a VQ-VAE trained over the last layer of hidden states. We also partition each layer’s prefix hidden states into contiguous, non-overlapping blocks of length 16 and train a VQ-VAE to acquire the code $\mathbf{B}_{k}^{\ell}$ where the block index $k\in\{1,\dots,12\}$ denotes the $k^{\text{th}}$ block from the end of the prefix (i.e., $k=1$ corresponds to the most recent 16 time steps before generation). In Figure 2 (middle), a sample block and a sample last hidden state are illustrated. We quantify the dependence between the codes of each computation block and the decision state by measuring nMI between $\mathbf{B}_{k}^{\ell}$ and $Z_{\scriptstyle T+\tau}^{\scriptstyle L}$} across all layers $\ell$ and blocks $k$, as defined in Eq. 2, with $\mathcal{I}_{\max}=\max\{\mathcal{I}(\mathbf{B}_{k}^{\ell};Z^{L}_{T+n})\}_{\ell,k,n}$. Heatmaps in Figure 4 (left and middle) show the results for $\tau$= 0 and 1 (see App. B.5.1 for $\tau>1$).
We observe that last layer computations retain the most information about the decision state of both immediate and future tokens which is aligned with the results of Pal et al. (2023) and consistent with the design choice of assigning longer attention spans to higher layers (Sukhbaatar et al., 2019; Beltagy et al., 2020). Furthermore, along the block axis, we observe a clear recency effect across all layers: the codes of the most recent blocks, i.e., final time steps of the prefix, computations ($Z_{1}$), exhibit the highest nMI (Eq. 2) with the decision state of the generated tokens, and the nMI decays over earlier blocks in the prefix. Although analysis along both axes indicates that the LM primarily relies on the most recent computations, we still observe appreciable nMI in lower layers (small $\ell$) and earlier blocks (large $k$). This suggests that LMs retain information from earlier parts of the prefix when generating new tokens, instead of relying only on the most recent computations.